What is the domain of Q(x)= x+8/x^2+36x
Answer by solver91311(24713)



That depends a great deal on how much of your function that follows the "/" is in the denominator of the rational part of the function, to wit:



has a different domain than



and both have a different domain than:



Repost your question using parentheses to eliminate ambiguities, that is either:

Q(x) = x + (8/x^2) + 36x

Q(x) = x + 8/(x^2 + 36x)

Q(x) = x + 8/x^(2 + 36x)

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism